OpenStudy (anonymous):

Okay, this problem is a Quadratic Function in Vertex Form problem: y= -(x + 3)^2 + 5

OpenStudy (anonymous):

you need to find values of x ?

OpenStudy (anonymous):

How would I do that?

OpenStudy (anonymous):

- ( x + 3 ) ^2 can you expand this ?

OpenStudy (plainntall):

What is the question? y=(x-k)^2 + k is the vertex form of a parabola

OpenStudy (plainntall):

sorry y=(x-h)^2 + k

OpenStudy (anonymous):

So just plugin?

OpenStudy (plainntall):

If you need to solve for x just expand like antoni7 said. If you need the vertex that is a different problem.

OpenStudy (anonymous):

How do I expand it?

OpenStudy (plainntall):

y= -(x + 3)^2 + 5 you need to multiply x+5 by itself y=-(x+3)(x+3)+5=?

OpenStudy (anonymous):

So that would equal 11

OpenStudy (dejavo):

x=-3

OpenStudy (plainntall):

no, if you want to find the values of x so that y=0 you have to solve for x what is (x+3)(x+3)? use FOIL method (First Outer Inner Last)

OpenStudy (plainntall):

Are you sure you have to solve for x? Usually you don't put it in vertex form for that. usually it has to do with graphing the function or finding the vertex.

OpenStudy (anonymous):

I don't think so, because in class we used y+ a (x-h)^2 + K to do Vertex, Axis of Symmetry, and then the graph.

OpenStudy (plainntall):

Okay so when you use the foil method to square x+3 you go first outer inner last x*x+3x+3x+3*3 = ?